OpenStudy (anonymous):

what is the value of the discriminant of the polynomial 4x^2+7x+4?

OpenStudy (baldymcgee6):

The discriminant is referring to b^2-4ac When a function is in the form of y = ax^2+bx+c

OpenStudy (baldymcgee6):

So in this case: 4x^2+7x+4 b^2-4ac (7)^2 - (4*4*4)

OpenStudy (anonymous):

thank you

OpenStudy (maheshmeghwal9):

\[ax^2+bx+c=0\]\[a, b \in \mathbb{R} \space ; \space \space a \neq 0\]\[\text{Discriminant}=D=\color{blue}{\Delta}=\color{red}{b^2-4ac.}\]